# Harmonic Loss Trains Interpretable AI Models

### v1

@@ 5 MNIST Experiments
Figure 5: Visualization of model weights trained for MNIST. Yellow cells show values less than $0.01$. Both models achieved $\approx 92.5\%$ test accuracy.
For vision tasks, convolutional neural networks are shown to be (at least somewhat) interpretable by demonstrating “edge detectors”, “wheel detectors”, etc. (Olah et al., 2020b). However, fully connected networks do not appear to be interpretable (e.g., see Figure 5 top). In this section, we demonstrate that the harmonic loss can lead to a more interpretable network for the MNIST dataset.
As a proof of concept, we choose our setup to be as simple as possible. Thus, we compare 1-layer neural networks trained using cross-entropy loss and harmonic loss. The input images are first flattened and passed through a $784\times 10$ linear layer to obtain the logits. The models were trained with a batch size of 64, a learning rate of 0.001, and for 10 epochs, achieving a 92.50% test accuracy for cross-entropy loss and 92.49% test accuracy for harmonic loss. We now analyze the weights learned by the neural network.
Figure 5 verifies that the weights in the model trained with harmonic loss are highly interpretable. Consistent with its core principle, we observe that the weights align with the class centers, which, in this case, correspond to images representing each number. Additionally, most peripheral pixels have weights that are almost exactly zero, in contrast to the model trained with cross-entropy loss. The latter, by design, lacks an incentive to reduce the irrelevant background weights to zero.
@@ 6 GPT2 Experiments
Figure 6: GPT2 experiments, trained on OpenWebText for 10k steps. (Top left) loss curves. Harmonic GPT achieves a slightly lower loss compared to the standard GPT. (Top right) accumulated distribution function with respect to parallelogram loss, for twelve function-vector tasks. The Harmonic GPT consistently shows lower parallelogram losses (i.e., better parallelograms). (Bottom) Parallelograms (1st and 2nd principal component) with quality ranked in descending order from left to right. The Harmonic GPT tends to produce parallelograms that are more ‘rectangular’, while standard GPT produces flat ‘parallelograms’.
Many mechanistic interpretability works have been dedicated to understanding large language models. For example, probing and attribution methods are good post hoc analysis tools. Despite their (partial) success, these tools are not creating interpretable models in the first place but are trying to find needles in the haystack. We argue that it would be nicer if we could pre-train the language models to be more interpretable. By using harmonic loss in training, we can produce a language model that can “grow” crystal-like representations, while having comparable performance with a standard one (trained with the cross-entropy loss).
@@
To testify the interpretability of the learned embeddings, we take twelve function-vector tasks from (Todd et al., 2023). Each dataset contains many input-output pairs that have a certain relation. For example, the “present-past” dataset contains pairs like: jump-jumped, fasten-fastened, win-won, etc. To construct parallelograms, we can draw two different pairs from the dataset, obtaining quadruples like (jump, jumped, fasten, fastened) which are expected to form parallelograms. Each word is tokenized into tokens; if multiple tokens are obtained, we use the last token. We project token embeddings onto the first two principal components. The quadruple $(i,j,m,n)$ has 2D PC embeddings $(\bm{E}_{i},\bm{E}_{j},\bm{E}_{m},\bm{E}_{n})$; we define the parallelogram loss $l_{\rm para}$ to be
@@
where $\sigma=\sqrt{\frac{1}{V}\sum_{k=1}^{V}\|\bm{E}_{k}\|^{2}}$ is a scale factor that normalizes the loss ($\bm{E}_{k}\to a\bm{E}_{k}$ leaves $l_{\rm para}$ invariant). We obtain 10000 quadruples, measuring the parallelogram qualities by computing their parallelogram losses. We plot their cumulated distribution function in Figure 6 top right: for every task, the harmonic GPT produces lower parallelogram loss (better parallelograms) than standard GPT. We show the parallelograms obtained in the present-past task in Figure 6 bottom. The parallelograms are ranked with quality in descending order from left to right. The harmonic GPT tends to produce visually appealing parallelograms that are more ‘rectangular’, while standard GPT produces flat ‘parallelograms’.
@@ 7 Conclusions
In this paper, we introduced harmonic loss as an alternative to the standard cross-entropy loss for training neural networks and large language models (LLMs). We found that models trained with harmonic loss outperform standard models by: (a) reducing grokking, (b) requiring less data for generalization, and (c) improving interpretability. We also compared a GPT-2 model trained with harmonic loss to the standard GPT-2, illustrating that the harmonic loss-trained model develops more interpretable representations. Further study is needed to explore the scalability and applicability of our findings to even larger models.
Impact Statement
This paper presents work whose goal is to enhance the interpretability of machine learning systems. Our harmonic loss approach enables a deeper understanding of model behavior, thereby improving the trustworthiness of machine learning systems.

### v2

@@ 4 MNIST Experiments
For vision tasks, convolutional neural networks are shown to be (at least somewhat) interpretable by demonstrating “edge detectors”, “wheel detectors”, etc. [9]. In this section, we show that the harmonic loss can lead to a more interpretable network for the MNIST dataset when it comes to training fully connected networks. As a proof of concept, we compare one-layer neural networks trained using cross-entropy loss and harmonic loss. The input images are first flattened and passed through a $784\times 10$ linear layer to obtain the logits. The models were trained with a batch size of 64, a learning rate of 0.001, and for 10 epochs, achieving a 92.50% test accuracy for cross-entropy loss and 92.49% test accuracy for harmonic loss.
Figure 4 shows that the harmonic model’s weights are more interpretable than those of the standard model. Consistent with its core principle, the harmonic model’s weights almost perfectly align with class centers (images of each number). They also assign near-zero values to peripheral pixels, unlike the model trained with cross-entropy loss, which lacks an incentive to reduce irrelevant background weights to exactly zero.
@@ 5 GPT2 Experiments
Figure 5: GPT2 experiments: (Top left) loss curves. Harmonic GPT achieves a slightly lower loss compared to standard GPT. (Top right) cumulative distribution function with respect to parallelogram loss, for twelve function-vector tasks. Harmonic GPT consistently shows lower parallelogram losses (i.e., better parallelograms). (Bottom) Parallelograms (1st and 2nd principal component) with quality ranked in descending order from left to right. Harmonic GPT tends to produce parallelograms that are more ‘rectangular’, while standard GPT produces flat ‘parallelograms’.
Many mechanistic interpretability works have been dedicated to understanding large language models. For example, probing and attribution methods are good post hoc analysis tools. Despite their (partial) success, these tools are not creating interpretable models in the first place but are trying to find needles in the haystack. We argue that it would be nicer if we could pre-train the language models to be more interpretable. By using harmonic loss in training, we can produce a language model that can “grow" crystal-like representations, while having comparable performance with a standard one (trained with the cross-entropy loss).
@@
To testify the interpretability of the learned embeddings, we take twelve function-vector tasks from [10]. Each dataset contains many input-output pairs that have a certain relation. For example, the “present-past" dataset contains pairs like: jump-jumped, fasten-fastened, win-won, etc. To construct parallelograms, we can draw two different pairs from the dataset, obtaining quadruples like (jump, jumped, fasten, fastened) which are expected to form parallelograms. Each word is tokenized into tokens; if multiple tokens are obtained, we use the last token. We project token embeddings onto the first two principal components. The quadruple $(i,j,m,n)$ has 2D PC embeddings $(\bm{E}_{i},\bm{E}_{j},\bm{E}_{m},\bm{E}_{n})$; we define the parallelogram loss $l_{\rm para}$ to be
@@
where $\sigma=\sqrt{\frac{1}{V}\sum_{k=1}^{V}\|\bm{E}_{k}\|^{2}}$ is a scale factor that normalizes the loss ($\bm{E}_{k}\to a\bm{E}_{k}$ leaves $l_{\rm para}$ invariant). We obtain 10000 quadruples, measuring the parallelogram qualities by computing their parallelogram losses. We plot their cumulative distribution function in Figure 5 in the top right: for every task, the harmonic GPT produces lower parallelogram loss (better parallelograms) than standard GPT. We show the parallelograms obtained in the present-past task in Figure 5 bottom. The parallelograms are ranked with quality in descending order from left to right. The harmonic GPT tends to produce visually appealing parallelograms that are more ‘rectangular’, while standard GPT produces flat ‘parallelograms’. Discussion about internal representations is included in Appendix C.
Figure 6: Learned embeddings on the lattice and modular addition tasks. Each pane shows the $5{\times}5$ class embeddings after training (numbers denote class IDs). Columns vary random seeds; the four left columns are for in-context learning, and the four right columns are for modular addition task. Rows correspond to loss functions: (top) full harmonic loss ($\ell_{2}$ logits + HarMax), (2nd) $\ell_{2}$ logits + SoftMax, (3rd) dot-product logits + HarMax, (bottom) standard cross-entropy layer. Here, we see that only $\ell_{2}$ distance paired with HarMax successfully recovers both the lattice and circular structure.
6 Ablation Experiments
Harmonic loss makes two major modifications to the standard cross-entropy loss: (i) compute logits via $\ell_{2}$ distances, and (ii) use HarMax function as shown in Eq. (2). To tease apart their individual contributions, we perform a set of targeted ablations in which one component is replaced at a time while the remainder of the training pipeline is left unchanged. Specifically, we train MLP models on the in-context learning and modular addition tasks with the ablated loss functions.
Results are shown in Figure 6. In in-context learning tasks, we observe that including either HarMax or $\ell_{2}$ logits alone is sufficient to replicate the full performance of Harmonic Loss. In contrast, for modular addition tasks, both HarMax and $\ell_{2}$ logits are essential to achieve the full performance. While incorporating only one component enhances the quality of the circular representation, the explained variance remains significantly below 100%. Overall, both HarMax and $\ell_{2}$ logits play critical roles in improving interpretability of the representations.
7 Related Works
Representations and Mechanistic Interpretability: Numerous studies have shown that LLMs can form conceptual representations across spatial [11], temporal [12], and color domains [13]. The structure of such representations includes one-dimensional concepts [11, 14, 15, 16], as well as multi-dimensional representations such as lattices [17, 18, 19] and circles [20, 21]. While the structure of these representations correlates with certain geometric patterns, significant unexplained variance frequently remains, necessitating efforts to improve the interpretability of neural network representations.
Loss Functions: Previous research has shown that loss functions can influence how a model learns to represent data, affecting its abilities in unique ways [22, 23, 24, 25, 26, 27, 28]. We refer readers to [29] and [30] for a comprehensive survey of different loss functions used in machine learning. Our harmonic loss offers an alternative supervisory signal in standard supervised learning by (a) replacing the usual SoftMax normalization with a scale-invariant HarMax function and (b) computing logits via Euclidean distance rather than a dot product. While it bears resemblance to contrastive loss—since both encourage maximal separation between different classes by using Euclidean distance as a metric—contrastive learning methods are not inherently supervised: they typically append a cross-entropy layer to generate logits, thus reintroducing SoftMax (and its drawbacks). We also show in Section 6 that using Euclidean distance alone is insufficient to fully replicate harmonic loss’s capabilities. Furthermore, directly leveraging Euclidean distance-based supervised learning has been relatively underexplored in language modeling outside of simple tasks like sentence sentiment classification [31]. We present a more comprehensive comparison of harmonic loss with other loss functions in Appendix D.
@@ 8 Conclusions
In this paper, we introduced harmonic loss as an alternative to the standard cross-entropy loss for training neural networks and large language models (LLMs). We found that models trained with harmonic loss perform better than standard models by: (a) reducing grokking, (b) requiring less data for generalization, and (c) improving interpretability. We also compared a GPT-2 model trained with harmonic loss to the standard GPT-2, illustrating that the harmonic loss-trained model develops more interpretable representations. Further study is needed to explore the scalability and applicability of our findings to even larger models.
Acknowledgements
This work is supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/).
@@ Appendix B Identifying Coset Structure in Permutation Representations
Appendix C Analyzing GPT2 hidden representations
In Section 5, we have shown that GPT2 trained with the harmonic loss has nicer structures in its embeddings (i.e., parallelograms) than that trained with the standard cross-entropy loss. We now show that intermediate representations (output of Block 6) induced by the harmonic loss are also qualitatively different from those of the cross-entropy loss. In Figure 8, the harmonic loss produces more perfect parallelograms (spike around zero parallelogram loss) but also displays a heavier tail for the parallelogram loss. The heavy tail is due to the heavy-tailedness of the harmonic loss (power law), as opposed to the cross-entropy loss (exponential). It remains to be understood if such heavy-tailedness is a feature or a bug for the harmonic loss, but the more perfect parallelograms are probably a good thing, or this at least suggests that imposing the harmonic loss at the end of the network can have noticeable influences in the intermediate representations. In Figure 9, we also notice that for the Captalize dataset, the lowercase and uppercase words tend to overlap in the first two PCs with the harmonic loss, but not with the cross-entropy loss. This again suggests the qualitative difference between the harmonic loss and the cross-entropy loss.
Figure 8: Harmonic loss (harmonic) and cross-entropy loss (standard) induce qualitatively different representations in the intermediate layer 6 of GPT2. We show the distribution of parallelogram loss for the parallelogram dataset. Harmonic loss has more perfect parallelograms (spike close to zero loss) but demonstrates a heavier tail.
Figure 9: Visualization of layer 6 representations projected onto the first two principal components, for the capitalize dataset. The harmonic loss (bottom) tends to collapse corresponding lower-case and upper-case words, while the cross-entropy loss (top) places them at different locations.
Appendix D Comparison of Harmonic Loss to Alternative Loss Functions
We briefly contrast the harmonic layer ($\ell_{2}$ logits + HarMax) with three popular loss families. Throughout, let $\mathbf{x}$ be an example embedding, $\mathbf{w}_{y}$ the weight of the correct class $y$, and $\mathbf{w}_{i}$ those of incorrect classes.
(a) Contrastive / InfoNCE.
A generic form is
$\mathcal{L}_{\text{contr}}\!=\!-\log\frac{\exp\!\bigl{(}s(\mathbf{x},\mathbf{x}^{+})/\tau\bigr{)}}{\exp\!\bigl{(}s(\mathbf{x},\mathbf{x}^{+})/\tau\bigr{)}+\sum_{i}\exp\!\bigl{(}s(\mathbf{x},\mathbf{x}_{i}^{-})/\tau\bigr{)}}.$
It enforces only relative ordering $s(\mathbf{x},\mathbf{x}^{+})>s(\mathbf{x},\mathbf{x}^{-})+m$, so entire constellations can drift or rotate. In contrast, harmonic loss pulls every example directly toward a fixed class anchor $\mathbf{w}_{y}$ and repels it from all others, yielding a stable, globally referenced geometry.
(b) Margin-based SoftMax.
Large-margin variants add a fixed gap $\Delta$ to every class boundary, $s(\mathbf{x},\mathbf{w}_{y})\geq s(\mathbf{x},\mathbf{w}_{i})+\Delta$. Because $\Delta$ is global, semantically close classes (e.g. dog vs. cat) are forced as far apart as unrelated ones (dog vs. airplane). Harmonic loss adapts separation dynamically: $p_{i}\propto\|\mathbf{x}-\mathbf{w}_{i}\|^{-n},$ so related concepts converge while unrelated ones diverge, yielding meaningful hierarchies (e.g. the Family-tree task).
(c) Spherical / cosine losses.
These constrain embeddings to the unit hypersphere and optimise angular margins: $\mathcal{L}_{\text{sph}}=-\log\frac{e^{s\,\cos\theta_{y}}}{\sum_{i}e^{s\,\cos\theta_{i}}}.$ While scale-invariant in angular space, they ignore absolute Euclidean proximity; our tasks (lattice, modular-add) benefit from the latter, explaining the poorer alignment of spherical loss.
We also run some experiments contrasting harmonic loss with loss (a) contrastive loss and (c) spherical loss for the in-context learning and modular addition tasks. Results for MLP and Transformer models are in Figure 10 and Figure 11, respectively.
Figure 10: Results for MLP models. Rows show harmonic, DotProd+HarMax, $\ell_{2}$ +SoftMax, and standard losses (top to bottom). Harmonic loss achieves the best reconstruction across seeds.
Figure 11: Results for Transformer models. Same ordering as Fig. 10.
Appendix E Sweeping HarMax Exponent Value
Figure 12: Effect of the harmonic exponent $n$ on lattice in-context learning. We sweep $n\in\{1,\dots,10\}$. Columns 1–4: Harmonic–MLP, columns 5–8: Harmonic Transformer. The learned $5\times 5$ lattice is remarkably stable; $n{=}1$ already provides crisp and interpretable geometry.
Figure 13: Effect of the harmonic exponent $n$ on modular addition. Columns 1–4: Harmonic–MLP, columns 5–8: Harmonic Transformer. MLPs remain stable across seeds, whereas Transformers are more sensitive yet form tighter circles at higher $n$; $n{=}1$ works well for MLPs, while a larger $n$ may benefit Transformers.
We perform experiments sweeping the HarMax exponent value for the in-context learning and modular addition tasks. Results are displayed in Figure 12 and Figure 13. We note that varying $n$ has minor impacts on lattice quality, with the default choice $n\!=\!1$ having the highest explained variances. Based on the modular addition task, our overall takeaway is that MLPs prefer the default $n{=}1$, while explained variance and circular structure for Transformer representations may improve with a slightly larger exponent.
Appendix F Full Results on Algorithmic Datasets
Fig. 14 shows the full results on algorithmic datasets.
Figure 14: (a) Cumulative explained variance vs. principal components (mean over 20 seeds). Harmonic representations are more compact than standard counterparts. (b) Test Accuracy as a function of Train Fraction (fixed seed). Harmonic models generalize faster with less data than standard counterparts. (c) Epochs to Test Acc $>$ 0.9 vs. Epochs to Train Acc $>$ 0.9 for 20 consecutive epochs. $y=x$ line represents no grokking, where train and test accuracy improve simultaneously. Points closer to the y-axis indicate a greater degree of grokking. Results from 20 different random seeds are plotted, and the runs that were not able to achieve 90% accuracy were omitted.
Appendix G Properties of Harmonic Loss: Proofs
Theorem 1 (Finite Convergence of Harmonic Loss).
Consider a classification model with $K$ classes and weight vectors $w_{1},\dots,w_{K}\in\mathbb{R}^{d}$ (no bias). Let $\{(x_{i},y_{i})\}_{i=1}^{n}$ be the training set, with $y_{i}\in\{1,\dots,K\}$. The cross-entropy loss is given by
$L_{\mathrm{CE}}(W)=-\sum_{i=1}^{n}\ln\frac{\exp(w_{y_{i}}\cdot x_{i})}{\sum_{j=1}^{K}\exp(w_{j}\cdot x_{i})}.$
The harmonic loss (with exponent $\beta>0$) is given by
$L_{\mathrm{H}}(W)=-\sum_{i=1}^{n}\ln\frac{\|x_{i}-w_{y_{i}}\|^{-\beta}}{\sum_{j=1}^{K}\|x_{i}-w_{j}\|^{-\beta}}\,.$
If the training data is linearly separable (i.e. there exists $W$ such that for all $i$, $w_{y_{i}}\cdot x_{i}>w_{j}\cdot x_{i}$ for $j\neq y_{i}$), then:
  • $L_{\mathrm{CE}}(W)$ has no finite minimum. In fact, for any weight matrix $W$ that classifies all training points correctly, one can decrease $L_{\mathrm{CE}}$ further by scaling $W$ to larger norm. Thus the infimum of $L_{\mathrm{CE}}$ is 00 but it is approached only as $\|W\|\to\infty$.
  • $L_{\mathrm{H}}(W)$ attains a (global) minimum at some finite $W$. Once the weights are large enough to classify all training points correctly (i.e. $\|x_{i}-w_{y_{i}}\|<\min_{j\neq y_{i}}\|x_{i}-w_{j}\|$ for all $i$), increasing the norm of $W$ does not reduce $L_{\mathrm{H}}$. In particular, $L_{\mathrm{H}}$ is scale-invariant: scaling all $w_{k}$ and all $x_{i}$ by a common factor leaves the loss unchanged. Consequently, $L_{\mathrm{H}}$ has a finite global minimizer.
Proof.
For the cross-entropy loss $L_{\mathrm{CE}}$, suppose $W$ classifies all training examples correctly. Then for each $i$, $w_{y_{i}}\cdot x_{i}>\max_{j\neq y_{i}}w_{j}\cdot x_{i}$. Consider scaling $W$ by a factor $t>1$: replace each $w_{k}$ with $tw_{k}$. Then $w_{y_{i}}\cdot x_{i}$ and $w_{j}\cdot x_{i}$ are both multiplied by $t$. The SoftMax probability of the true class $y_{i}$ becomes
$P_{W}(y_{i}|x_{i})=\frac{\exp(w_{y_{i}}\cdot x_{i})}{\sum_{j}\exp(w_{j}\cdot x_{i})}.$
Under scaling $tW$, this becomes
$P_{tW}(y_{i}|x_{i})=\frac{\exp(t\,w_{y_{i}}\cdot x_{i})}{\sum_{j}\exp(t\,w_{j}\cdot x_{i})}.$
Since $w_{y_{i}}\cdot x_{i}$ is the largest logit for sample $i$, as $t\to\infty$ we have $P_{tW}(y_{i}|x_{i})\to 1$ and thus $-\ln P_{tW}(y_{i}|x_{i})\to 0$. This holds for all $i$, so $L_{\mathrm{CE}}(tW)\to 0$ as $t\to\infty$. Therefore, no finite $W$ minimizes $L_{\mathrm{CE}}$; the infimum 00 is approached only in the limit $\|W\|\to\infty$.
For $L_{\mathrm{H}}$, once $W$ is such that each training point is correctly classified by its nearest prototype (i.e. $\|x_{i}-w_{y_{i}}\|<\|x_{i}-w_{j}\|$ for all $j\neq y_{i}$), increasing the norms $\|w_{k}\|$ further will not improve the loss. In fact, if every $x_{i}$ is closer to its correct $w_{y_{i}}$ than to any other $w_{j}$, then the harmonic probabilities
$P_{W}(y_{i}|x_{i})=\frac{\|x_{i}-w_{y_{i}}\|^{-\beta}}{\sum_{j=1}^{K}\|x_{i}-w_{j}\|^{-\beta}}$
remain unchanged under a uniform scaling: if we replace $x_{i}$ by $cx_{i}$ and $w_{k}$ by $cw_{k}$, then $\|cx_{i}-cw_{k}\|=c\,\|x_{i}-w_{k}\|$, so the scaling factors cancel. Therefore, once correct classification is achieved, no further reduction in loss is obtained by increasing $\|W\|$, and $L_{\mathrm{H}}$ achieves its minimum at finite $W$. ∎
Theorem 2 (PAC-Bayesian Generalization Bound of Harmonic Loss).
Assume all training examples lie within a ball of radius $R$ in input space, i.e. $\|x_{i}\|\leq R$ for all $i$. Suppose a weight matrix $W$ achieves a distance margin of $\gamma>0$ on the training set, meaning that for every training sample $(x_{i},y_{i})$ and any other class $j\neq y_{i}$,
$\|x_{i}-w_{y_{i}}\|+\gamma\leq\|x_{i}-w_{j}\|.$
Then, with probability at least $1-\delta$, the generalization (test) error of the harmonic classifier satisfies
$\Pr_{(x,y)\sim D}\big{[}h_{W}(x)\neq y\big{]}\;\leq\;\mathcal{O}\left(\frac{R\,\|W\|}{\gamma\sqrt{n}}+\sqrt{\frac{\ln(1/\delta)}{n}}\right),$
where $h_{W}(x)$ denotes the predicted class and $n$ is the number of training samples.
In particular, $\|W\|$ is finite for harmonic loss (by Theorem 1), and typically much smaller than the weight norm of the solution obtained with cross-entropy loss. Thus, the harmonic classifier has a tighter generalization bound.
Proof.
Applying the standard PAC-Bayes margin bounds (see e.g. [32]), one obtains that with probability at least $1-\delta$,
$\Pr(h_{W}(x)\neq y)\leq\mathcal{O}\left(\frac{R\,\|W\|}{\gamma\sqrt{n}}+\sqrt{\frac{\ln(1/\delta)}{n}}\right).$
Since the harmonic loss yields a solution with finite $\|W\|$, the bound is finite. In contrast, the cross-entropy solution would have $\|W\|\to\infty$ even when achieving zero training error, rendering a similar bound meaningless. ∎
Theorem 3 (Interpretable Representations of Harmonic Loss).
At a critical point (in particular, a global minimum) of the harmonic loss, each weight vector $w_{k}$ becomes an interpretable class center for class $k$. Specifically, the stationarity condition implies
$w_{k}\;=\;\sum_{i:y_{i}=k}\alpha_{i}\,x_{i}\qquad\text{with }\alpha_{i}\geq 0,\;\sum_{i:y_{i}=k}\alpha_{i}=1,$
i.e. $w_{k}$ is a convex combination of the training examples of class $k$. Consequently, $w_{k}$ represents the center point of its class, leading to more interpretable representations compared to cross-entropy loss.
Proof.
Differentiate the harmonic loss with respect to $w_{k}$. For simplicity, denote
$p_{i}^{k}=\frac{\|x_{i}-w_{k}\|^{-\beta}}{\sum_{j=1}^{K}\|x_{i}-w_{j}\|^{-\beta}}.$
For samples $x_{i}$ with $y_{i}=k$, the derivative takes the form
$\frac{\partial L_{\mathrm{H}}}{\partial w_{k}}=-\sum_{i:y_{i}=k}\frac{\beta}{\|x_{i}-w_{k}\|^{2}}(w_{k}-x_{i})\,p_{i}^{k}+\text{terms from }i\text{ with }y_{i}\neq k.$
At a critical point, the total derivative vanishes. Rearranging the stationarity conditions (and noting that the repulsive forces from other classes tend to balance out overall on average due to long distance) yields
$w_{k}=\frac{\sum_{i:y_{i}=k}\frac{1}{\|x_{i}-w_{k}\|^{2}}x_{i}+\sum_{j:y_{j}\neq k}\frac{1}{\|x_{j}-w_{k}\|^{2}}x_{j}}{\sum_{i:y_{i}=k}\frac{1}{\|x_{i}-w_{k}\|^{2}}+\sum_{j:y_{j}\neq k}\frac{1}{\|x_{j}-w_{k}\|^{2}}}.$
Since $w_{k}$ is closer to class-$k$ examples than to others, the weights $\frac{1}{\|x_{i}-w_{k}\|^{2}}$ for $i$ with $y_{i}=k$ dominate the sum. Define
$\alpha_{i}=\frac{\frac{1}{\|x_{i}-w_{k}\|^{2}}}{\sum_{i:y_{i}=k}\frac{1}{\|x_{i}-w_{k}\|^{2}}+\sum_{j:y_{j}\neq k}\frac{1}{\|x_{j}-w_{k}\|^{2}}}.$
Then $w_{k}$ can be written as a convex combination
$w_{k}=\sum_{i:y_{i}=k}\alpha_{i}\,x_{i}+\sum_{j:y_{j}\neq k}\alpha_{j}\,x_{j}\,.$
In many practical settings, the contribution from $x_{j}$ with $y_{j}\neq k$ is negligible, so $w_{k}$ is nearly a convex combination solely of class-$k$ samples. By construction, $\alpha_{i}\geq 0$ and the weights sum to 1. This shows that $w_{k}$ is an interpretable vector representing its class center. In contrast, for cross-entropy loss the stationary condition does not yield a similar expression for $w_{k}$ as a combination of data points. ∎
Remark: Under cross-entropy loss, the weight vectors usually end up pointing to the average direction of class elements, due to its use of the dot product. However, they do not have a closed-form formula like the harmonic loss above, and the weight vectors are not linear combinations of all class feature directions. We believe that enforcing such linear combination structure plays a crucial role in enhancing interpretability – it directly aligns with the Linear Representation Hypothesis [33], and natively supports compositional generalization.
Appendix H Additional Benchmark Results
H.1 ImageNet
ImageNet [34] is a large-scale visual dataset commonly used in object recognition research. We compare the performance of standard cross entropy loss and harmonic loss on ImageNet. We trained ResNet-50 with AutoAugment data augmentation method for 90 epochs, starting with a learning rate of 0.1, which was reduced by a factor of 10 at epochs 10, 30, 60, and 80. The training results are presented in Table 1. We have also implemented our own cross-entropy training pipeline, and compared them with existing results in [35]. In our implementation, the harmonic model modestly outperformed the standard model.
Table 1: Validation accuracy on ImageNet using different loss functions.
Loss Top-1 Val Acc Top-5 Val Acc
Cross-Entropy (Ours) 74.17% 91.88%
Harmonic Loss (Ours) 75.08% 92.12%
Cross-Entropy [35] 77.6% 95.3%
Supervised Contrastive Loss [35] 78.7% 94.3%
Table 2: Probing F1 score on SST-2 and CoLA datasets.
Model SST2 (Layer 0) CoLA (Layer 0) SST2 (Layer 6) CoLA (Layer 6)
Cross-Entropy 76.2 $\pm$ 1.5% 73.9 $\pm$1.0 % 79.9 $\pm$ 1.3 % 78.2 $\pm$ 1.7%
Harmonic 77.9 $\pm$ 1.1% 74.3 $\pm$ 1.0 % 79.9 $\pm$ 1.7 % 77.1 $\pm$ 4.2%
H.2 SST2 and GLUE
We also compare the standard GPT2 and harmonic GPT2 with the GLUE benchmark below. We evaluate two tasks, COLA (linguistic acceptibility) [36] and SST2 (sentence sentiment classification) [37]. We train a 1-layer MLP probe with hidden dimension 16 that takes the model’s residual stream representation as an input, and outputs the label. Table 2 shows the F1 score of the probe on validation dataset.